# IPFormer: Visual 3D Panoptic Scene Completion with Context-Adaptive Instance Proposals

### v2

@@ 3.4 Encoding
To encode semantic information, we propagate the voxel proposals $\mathbf{V}_{P}$ through a 3D Local and Global Encoder (LGE) based on CGFormer cgformer . This encoder enhances the semantic and geometric discriminability of the scene volume by integrating TPV huang2023tpvformer and voxel representations, effectively capturing global and local contextual features, respectively. For an in-depth discussion, we direct readers to cgformer . The 3D LGE $\Phi_{\text{LGE}}$ transforms voxel proposals $\mathbf{V}_{P}$ into semantic voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$, formulated as $\mathbf{V}_{F}=\Phi_{\text{LGE}}(\mathbf{V}_{P})$. Building on these semantically encoded voxel features, we apply cross-attention (CA) to instance proposals $\mathbf{I}_{P}\in\mathbb{R}^{K\times C}$ (as queries) and semantic voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$ (as keys and values), producing updated instance features $\tilde{\mathbf{I}}_{P}^{\text{CA}}$. Specifically, the updated instance features are computed as:
@@ 3.5 Decoding
In the second training stage, a lightweight panoptic head decodes instance features $\mathbf{I}_{F}$ to semantic class logits $\mathbf{I}_{L}\in\mathbb{R}^{K\times L}$, where $K$ is the number of instances and $L$ is the number of semantic classes. These logits are further processed with voxel features $\mathbf{V}_{F}$ to perform instance-voxel alignment that eventually enables panoptic scene aggregation, inspired by PaSCo pasco . To elaborate, we align the instance features $\mathbf{I}_{F}$ with the voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$ to compute affinity scores $\mathbf{H}\in\mathbb{R}^{X\times Y\times Z\times K}$ by applying the dot product, resulting in $\mathbf{H}=\mathbf{V}_{F}\cdot\mathbf{I}_{F}^{\top}$. These scores are converted to predicted probabilities via $\mathbf{H}_{P}=\text{sigmoid}(\mathbf{H})$, where $\mathbf{H}_{P}$ denotes the model’s confidence in each voxel belonging to each instance. For panoptic aggregation pasco , voxel-wise logits $\mathbf{V}_{L}\in\mathbb{R}^{X\times Y\times Z\times L}$ are computed as $\mathbf{V}_{L}=\mathbf{H}_{P}\cdot\mathbf{I}_{L}$. It should be emphasized that voxel-wise logits do not represent the probability of belonging to a certain semantic class, but the confidence of belonging to one of the instances. Furthermore, semantic class IDs are assigned via argmax over $\mathbf{V}_{L}$ along $L$, and instance IDs via argmax over $\mathbf{H}_{P}^{\top}$ along $K$, yielding a complete panoptic scene. Note that the alignment process and the panoptic aggregation do not include learnable parameters.
@@ 3.6 Training Objective
where the weights for the cross-entropy loss, semantic scene-class affinity loss (SCAL), and geometric SCAL monoscene are set to $\lambda_{\text{ce}}=\lambda_{\text{sem}}=\lambda_{\text{geo}}=1$, respectively, and the weight for the depth loss cgformer is set to $\lambda_{\text{depth}}=10^{-4}$.
The second stage calculates losses mask-wise by comparing each ground-truth mask with the best matching predicted mask. Consistent with previous research pasco ; shi2024panossc , we find the best match by applying the Hungarian method kuhn1955hungarian to perform bipartite matching carion2020end , using an IoU threshold of $50\text{\%}$. This threshold follows the standard convention established in prior work on panoptic segmentation kirillov2019panoptic ; pasco , and the Hungarian algorithm ensures a globally optimal, permutation-invariant one-to-one assignment. After matching, the second stage optimizes
@@
where $\mathcal{L}_{\text{ce}}^{\text{mask}}$ represents mask-wise cross-entropy loss, $\mathcal{L}_{dice}$ is the dice loss of Milletari_2016_diceloss_vnet and $\mathcal{L}_{focal}$ is the focal loss based on Lin_2017_focalloss . In practice, we set the weights of these losses to $\lambda_{\text{mask}}=\lambda_{\text{dice}}=1$ and $\lambda_{\text{focal}}=40$, respectively, following PaSco pasco . Moreover, $\mathcal{L}_{\text{depth}}$ is defined identically to that in $\mathcal{L}_{\text{SSC}}$.
@@ 4.1 Quantitative Results
Dataset and Baselines.   We conduct our experiments by (i) in-domain training and testing on the SemanticKITTI SSC dataset semantickitti , and (ii) out-of-domain zero-shot generalization on the distinct SSCBench-KITTI360 li2024sscbench . The instance ground-truths for both datasets are provided by PaSCo pasco . Given the novelty of vision-based PSC, no established baselines currently exist for direct comparison. Furthermore, recall from Sec. 2 that PanoSSC shi2024panossc represents an existing vision-based PSC method using forward-facing imagery. This would make it a suitable baseline for comparison. However, this approach is trained on a non-public post-processed version of the SemanticKITTI dataset semantickitti , which, in addition to the absence of released source code, challenges direct comparison with our method (see Sec. A.6 of the technical appendix for further details). Consequently, we adapt the latest state-of-the-art vision-based SSC methods monoscene ; symphonies ; occformer ; cgformer and generate instance predictions by clustering Thing-classes from their outputs, in line with PaSCo pasco , which introduced the task of 3D PSC. In particular, to ensure fairness and methodological rigor, we apply DBSCAN ester1996_dbscan , the same Euclidean clustering algorithm used to construct the PSC ground truth in PaSCo. Furthermore, we utilize pre-trained checkpoints for the baselines, acquired from the official publicly available implementations. We provide further details in Sec. A.1 of the technical appendix.
@@
Out-of-Domain Zero-Shot Generalization Performance.   Trained on SemanticKITTI semantickitti , we cross-validate our method on the distinct SSCBench-KITTI360 dataset li2024sscbench . Results in Tab. 2 show that IPFormer demonstrates superior zero-shot generalization capability. Especially in comparison to its closest baseline, CGFormer+DBSCAN, which it outperforms across PSC and SSC metrics. These results highlight the generalization capability of IPFormer by leveraging out-of-domain image context to robustly initialize context-adaptive instance proposals, resulting in superior SSC and PSC performance.
Table 1: In-domain performance on SemanticKITTI val. set semantickitti . Best and second-best results are bold and underlined, respectively. Due to the absence of established baselines for vision-based PSC (see Sec. 4.1), we infer state-of-the-art SSC methods and apply DBSCAN ester1996_dbscan to retrieve instances.
@@
MonoScene monoscene + DBSCAN     $10.12$ $3.43$ $15.15$ $5.33$ $0.51$ $7.36$ $0.87$ $5.56$ $20.81$ $8.57$     $36.80$ $11.31$     $\underline{4.51}$
Symphonies symphonies + DBSCAN     $11.69$ $3.75$ $26.09$ $5.95$ $1.07$ $27.65$ $1.76$ $5.70$ $24.95$ $8.99$     $\underline{41.92}$ $15.02$     $4.54$
OccFormer occformer + DBSCAN     $11.25$ $4.32$ $24.19$ $6.69$ $0.68$ $21.47$ $1.15$ $6.96$ $26.16$ $10.73$     $36.43$ $13.51$     $4.70$
CGFormer cgformer + DBSCAN     $\underline{14.39}$ $\underline{6.16}$ $\mathbf{48.14}$ $\underline{9.48}$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $\underline{9.03}$ $\mathbf{50.82}$ $\underline{13.86}$     $\mathbf{45.98}$ $\mathbf{16.89}$     $4.70$
@@
Table 2: Out-of-domain zero-shot generalization performance of IPFormer and the closest baseline CGFormer+DBSCAN, by training on SemanticKITTI semantickitti and cross-validating on SSCBench-KITTI360 test set li2024sscbench . IPFormer demonstrates superior absolute and relative generalization performance across PSC and SSC metrics.
@@
CGFormer cgformer + DBSCAN     $14.39$ $6.16$ $\mathbf{48.14}$ $9.48$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $9.03$ $\mathbf{50.82}$ $13.86$     $\mathbf{45.98}$ $\mathbf{16.89}$
@@
CGFormer cgformer + DBSCAN     $8.44$ $1.08$ $17.82$ $1.87$ $\mathbf{0.53}$ $20.06$ $\mathbf{0.96}$ $1.48$ $16.19$ $2.54$     $28.11$ $9.44$
@@ 4.2 Ablation Study
Deep Supervision.  Inspired by PanoSSC shi2024panossc , we investigate deep supervision in the form of auxiliary losses, essentially supervising the attention maps of each layer during the encoding of the instance proposals (Tab. 5). Our findings show that guiding intermediate layers in such way degrades all metrics. Most notably, when deep supervision is not applied, SQ-Thing registers a substantial improvement of $102.61\text{\,}\%$.
@@
Figure 4: Qualitative results on the SemanticKITTI val. set semantickitti . Each top row illustrates purely semantic information, following the SSC color map. Each bottom row displays individual instances, with randomly assigned colors to facilitate differentiation. Note that we specifically show instances of the Thing-category for clarity.

### v3

@@ 3.4 Encoding
To encode semantic information, we propagate the voxel proposals $\mathbf{V}_{P}$ through a 3D Local and Global Encoder (LGE) based on CGFormer Yu et al. (2024). This encoder enhances the semantic and geometric discriminability of the scene volume by integrating TPV Huang et al. (2023) and voxel representations, effectively capturing global and local contextual features, respectively. For an in-depth discussion, we direct readers to Yu et al. (2024). The 3D LGE $\Phi_{\text{LGE}}$ transforms voxel proposals $\mathbf{V}_{P}$ into semantic voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$, formulated as $\mathbf{V}_{F}=\Phi_{\text{LGE}}(\mathbf{V}_{P})$. Building on these semantically encoded voxel features, we apply cross-attention (CA) to instance proposals $\mathbf{I}_{P}\in\mathbb{R}^{K\times C}$ (as queries) and semantic voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$ (as keys and values), producing updated instance features $\tilde{\mathbf{I}}_{P}^{\text{CA}}$. Specifically, the updated instance features are computed as:
@@ 3.5 Decoding
In the second training stage, a lightweight panoptic head decodes instance features $\mathbf{I}_{F}$ to semantic class logits $\mathbf{I}_{L}\in\mathbb{R}^{K\times L}$, where $K$ is the number of instances and $L$ is the number of semantic classes. These logits are further processed with voxel features $\mathbf{V}_{F}$ to perform instance-voxel alignment that eventually enables panoptic scene aggregation, inspired by PaSCo Cao et al. (2024). To elaborate, we align the instance features $\mathbf{I}_{F}$ with the voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$ to compute affinity scores $\mathbf{H}\in\mathbb{R}^{X\times Y\times Z\times K}$ by applying the dot product, resulting in $\mathbf{H}=\mathbf{V}_{F}\cdot\mathbf{I}_{F}^{\top}$. These scores are converted to predicted probabilities via $\mathbf{H}_{P}=\text{sigmoid}(\mathbf{H})$, where $\mathbf{H}_{P}$ denotes the model’s confidence in each voxel belonging to each instance. For panoptic aggregation Cao et al. (2024), voxel-wise logits $\mathbf{V}_{L}\in\mathbb{R}^{X\times Y\times Z\times L}$ are computed as $\mathbf{V}_{L}=\mathbf{H}_{P}\cdot\mathbf{I}_{L}$. It should be emphasized that voxel-wise logits do not represent the probability of belonging to a certain semantic class, but the confidence of belonging to one of the instances. Furthermore, semantic class IDs are assigned via argmax over $\mathbf{V}_{L}$ along $L$, and instance IDs via argmax over $\mathbf{H}_{P}^{\top}$ along $K$, yielding a complete panoptic scene. Note that the alignment process and the panoptic aggregation do not include learnable parameters.
@@ 3.6 Training Objective
where the weights for the cross-entropy loss, semantic scene-class affinity loss (SCAL), and geometric SCAL Cao and de Charette (2022) are set to $\lambda_{\text{ce}}=\lambda_{\text{sem}}=\lambda_{\text{geo}}=1$, respectively, and the weight for the depth loss Yu et al. (2024) is set to $\lambda_{\text{depth}}=10^{-4}$.
The second stage calculates losses mask-wise by comparing each ground-truth mask with the best matching predicted mask. Consistent with previous research Cao et al. (2024); Shi et al. (2024), we find the best match by applying the Hungarian method Kuhn (1955) to perform bipartite matching Carion et al. (2020), using an IoU threshold of $50\text{\%}$. This threshold follows the standard convention established in prior work on panoptic segmentation Kirillov et al. (2019b); Cao et al. (2024), and the Hungarian algorithm ensures a globally optimal, permutation-invariant one-to-one assignment. After matching, the second stage optimizes
@@
where $\mathcal{L}_{\text{ce}}^{\text{mask}}$ represents mask-wise cross-entropy loss, $\mathcal{L}_{dice}$ is the dice loss of Milletari et al. (2016) and $\mathcal{L}_{focal}$ is the focal loss based on Lin et al. (2017). In practice, we set the weights of these losses to $\lambda_{\text{mask}}=\lambda_{\text{dice}}=1$ and $\lambda_{\text{focal}}=40$, respectively, following PaSco Cao et al. (2024). Moreover, $\mathcal{L}_{\text{depth}}$ is defined identically to that in $\mathcal{L}_{\text{SSC}}$.
@@ 4.1 Quantitative Results
Dataset and Baselines.   We conduct our experiments by (i) in-domain training and testing on the SemanticKITTI SSC dataset Behley et al. (2019), and (ii) out-of-domain zero-shot generalization on the distinct SSCBench-KITTI360 Li et al. (2024). The instance ground-truths for both datasets are provided by PaSCo Cao et al. (2024). Given the novelty of vision-based PSC, no established baselines currently exist for direct comparison. Furthermore, recall from Sec. 2 that PanoSSC Shi et al. (2024) represents an existing vision-based PSC method using forward-facing imagery. This would make it a suitable baseline for comparison. However, this approach is trained on a non-public post-processed version of the SemanticKITTI dataset Behley et al. (2019), which, in addition to the absence of released source code, challenges direct comparison with our method (see Sec. A.6 of the technical appendix for further details). Consequently, we adapt the latest state-of-the-art vision-based SSC methods Cao and de Charette (2022); Jiang et al. (2024); Zhang et al. (2023); Yu et al. (2024) and generate instance predictions by clustering Thing-classes from their outputs, in line with PaSCo Cao et al. (2024), which introduced the task of 3D PSC. In particular, to ensure fairness and methodological rigor, we apply DBSCAN Ester et al. (1996), the same Euclidean clustering algorithm used to construct the PSC ground truth in PaSCo. Furthermore, we utilize pre-trained checkpoints for the baselines, acquired from the official publicly available implementations. We provide further details in Sec. A.1 of the technical appendix.
@@
Out-of-Domain Zero-Shot Generalization Performance.   Trained on SemanticKITTI Behley et al. (2019), we cross-validate our method on the distinct SSCBench-KITTI360 dataset Li et al. (2024). Results in Tab. 2 show that IPFormer demonstrates superior zero-shot generalization capability. Especially in comparison to its closest baseline, CGFormer+DBSCAN, which it outperforms across PSC and SSC metrics. These results highlight the generalization capability of IPFormer by leveraging out-of-domain image context to robustly initialize context-adaptive instance proposals, resulting in superior SSC and PSC performance.
Table 1: In-domain performance on SemanticKITTI val. set Behley et al. (2019). Best and second-best results are bold and underlined, respectively. Due to the absence of established baselines for vision-based PSC (see Sec. 4.1), we infer state-of-the-art SSC methods and apply DBSCAN Ester et al. (1996) to retrieve instances.
@@
MonoScene Cao and de Charette (2022) + DBSCAN     $10.12$ $3.43$ $15.15$ $5.33$ $0.51$ $7.36$ $0.87$ $5.56$ $20.81$ $8.57$     $36.80$ $11.31$     $\underline{4.51}$
Symphonies Jiang et al. (2024) + DBSCAN     $11.69$ $3.75$ $26.09$ $5.95$ $1.07$ $27.65$ $1.76$ $5.70$ $24.95$ $8.99$     $\underline{41.92}$ $15.02$     $4.54$
OccFormer Zhang et al. (2023) + DBSCAN     $11.25$ $4.32$ $24.19$ $6.69$ $0.68$ $21.47$ $1.15$ $6.96$ $26.16$ $10.73$     $36.43$ $13.51$     $4.70$
CGFormer Yu et al. (2024) + DBSCAN     $\underline{14.39}$ $\underline{6.16}$ $\mathbf{48.14}$ $\underline{9.48}$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $\underline{9.03}$ $\mathbf{50.82}$ $\underline{13.86}$     $\mathbf{45.98}$ $\mathbf{16.89}$     $4.70$
@@
Table 2: Out-of-domain zero-shot generalization performance of IPFormer and the closest baseline CGFormer+DBSCAN, by training on SemanticKITTI Behley et al. (2019) and cross-validating on SSCBench-KITTI360 test set Li et al. (2024). IPFormer demonstrates superior absolute and relative generalization performance across PSC and SSC metrics.
@@
CGFormer Yu et al. (2024) + DBSCAN     $14.39$ $6.16$ $\mathbf{48.14}$ $9.48$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $9.03$ $\mathbf{50.82}$ $13.86$     $\mathbf{45.98}$ $\mathbf{16.89}$
@@
CGFormer Yu et al. (2024) + DBSCAN     $8.44$ $1.08$ $17.82$ $1.87$ $\mathbf{0.53}$ $20.06$ $\mathbf{0.96}$ $1.48$ $16.19$ $2.54$     $28.11$ $9.44$
@@ 4.2 Ablation Study
Deep Supervision.  Inspired by PanoSSC Shi et al. (2024), we investigate deep supervision in the form of auxiliary losses, essentially supervising the attention maps of each layer during the encoding of the instance proposals (Tab. 5). Our findings show that guiding intermediate layers in such way degrades all metrics. Most notably, when deep supervision is not applied, SQ-Thing registers a substantial improvement of $102.61\text{\,}\%$.
@@
Figure 4: Qualitative results on the SemanticKITTI val. set Behley et al. (2019). Each top row illustrates purely semantic information, following the SSC color map. Each bottom row displays individual instances, with randomly assigned colors to facilitate differentiation. Note that we specifically show instances of the Thing-category for clarity.